OpenStudy (anonymous):

which of the following is the rational exponent of 6sqrt?

OpenStudy (anonymous):

f^6 f^1/6 f/6 6f

ganeshie8 (ganeshie8):

hint : \(\large \sqrt[n]{a} = a^{\frac{1}{n}}\)

OpenStudy (anonymous):

so b?

ganeshie8 (ganeshie8):

may be... your question is incomplete, check it once you have eaten some part of it :o

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

i agree with @ganeshie8 !

OpenStudy (anonymous):

that if i eat it ill find my answer?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

no

OpenStudy (cggurumanjunath):

check there is some mistake in the q ! i meant. @allie_bear22

OpenStudy (anonymous):

the expression is \[^6\sqrt{?}\]

OpenStudy (anonymous):

thats what it looks like

OpenStudy (anonymous):

@ganeshie8 @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):